# Rate-Distortion-Perception Tradeoff for Gaussian Vector Sources

### v1

@@ Appendix H Asymptotic Analysis for Perceptually Perfect Reconstruction
H-1 High-Distortion Compression
@@
$=$ $\frac{\lambda_{\ell}}{1+4\nu_{1}^{2}\lambda^{2}_{\ell}+O(\nu_{1}^{4})}$ (180)
$=$ $\lambda_{\ell}(1-4\nu_{1}^{2}\lambda^{2}_{\ell})+O(\nu_{1}^{4}).$ (181)
@@
$=$ $\frac{1}{2}\sum_{\ell=1}^{L}\log(1+4\nu_{1}^{2}\lambda_{\ell}^{2}+O(\nu_{1}^{4}))$ (184)
$=$ $\frac{1}{2}\sum_{\ell=1}^{L}4\nu_{1}^{2}\lambda_{\ell}^{2}+O(\nu_{1}^{4}),$ (185)
@@
To derive the expression for the water-level, we use (182) in (181) to get
$\gamma^{*}_{\ell}\left(2\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,0\right)=\lambda_{\ell}-\frac{\epsilon^{2}\lambda_{\ell}^{3}}{4\left(\sum_{\ell=1}^{L}\lambda_{\ell}^{2}\right)^{2}}+O(\epsilon^{3}),\quad\ell\in\{1,\ldots,L\}.$ (187)
H-2 Low-Distortion Compression
Let $D=\epsilon$ for some small $\epsilon>0$. Note that as $\epsilon\rightarrow 0$, we must have $\gamma_{\ell}^{*}\rightarrow 0$ by (177), and consequently $\nu_{1}\rightarrow\infty$ by (176). In this regime, we can approximate the water-levels in (176) as follows
@@ H-2 Low-Distortion Compression
$=$ $\frac{1}{2\nu_{1}}-\frac{1}{8\nu_{1}^{2}\lambda_{\ell}}+O\left(\frac{1}{\nu_{1}^{3}}\right).$ (189)
@@
$=$ $\frac{L}{2\nu_{1}}-\frac{1}{16\nu_{1}^{2}}\sum_{\ell=1}^{L}\frac{1}{\lambda_{\ell}}+O\left(\frac{1}{\nu_{1}^{3}}\right),$ (191)
@@
Substituting (192) into (189) shows that the water-levels in the low-distortion regime are given by
$\gamma_{\ell}^{*}(\epsilon,0)=\frac{\epsilon}{L}-\frac{\epsilon^{2}}{2L^{2}\lambda_{\ell}}+\frac{\epsilon^{2}}{4L^{3}}\sum_{\ell=1}^{L}\frac{1}{\lambda_{\ell}}+O(\epsilon^{3}),\quad\ell\in\{1,\ldots,L\}.$ (193)
@@
$R(\epsilon,0)$ $=$ $\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{\lambda_{\ell}}{\gamma^{*}_{\ell}(\epsilon,0)}$ (194)
$=$ $\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{L\lambda_{\ell}}{\epsilon}-\frac{1}{2}\sum\limits_{\ell=1}^{L}\log\left(1-\frac{\epsilon}{2L\lambda_{\ell}}+\frac{\epsilon}{4L^{2}}\sum_{\ell^{\prime}=1}^{L}\frac{1}{\lambda_{\ell^{\prime}}}+O(\epsilon^{2})\right)$ (195)
$=$ $\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{L\lambda_{\ell}}{\epsilon}-\frac{1}{2}\sum\limits_{\ell=1}^{L}\left(-\frac{\epsilon}{2L\lambda_{\ell}}+\frac{\epsilon}{4L^{2}}\sum_{\ell^{\prime}=1}^{L}\frac{1}{\lambda_{\ell^{\prime}}}\right)+O(\epsilon^{2})$ (196)

### v2

@@ Appendix H Asymptotic Analysis for Perceptually Perfect Reconstruction
H-A High-Distortion Compression
@@
$=$ $\frac{\lambda_{\ell}}{1+4\nu_{1}^{2}\lambda^{2}_{\ell}+O(\nu_{1}^{4})}$
$=$ $\lambda_{\ell}(1-4\nu_{1}^{2}\lambda^{2}_{\ell})+O(\nu_{1}^{4}).$
@@
$=$ $\frac{1}{2}\sum_{\ell=1}^{L}\log(1+4\nu_{1}^{2}\lambda_{\ell}^{2}+O(\nu_{1}^{4}))$
$=$ $\frac{1}{2}\sum_{\ell=1}^{L}4\nu_{1}^{2}\lambda_{\ell}^{2}+O(\nu_{1}^{4}).$
@@
To derive the expression for the water level, we use (185) in (184) to get
$\gamma^{*}_{\ell}\left(2\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,0\right)=\lambda_{\ell}-\frac{\epsilon^{2}\lambda_{\ell}^{3}}{4\left(\sum_{\ell=1}^{L}\lambda_{\ell}^{2}\right)^{2}}+O(\epsilon^{3}),$
$\hskip 142.26378pt\ell\in\{1,\ldots,L\}.$ (188)
H-B Low-Distortion Compression
Let $D=\epsilon$ for some small $\epsilon>0$. Note that as $\epsilon\rightarrow 0$, we must have $\gamma_{\ell}^{*}\rightarrow 0$ by (182), and consequently $\nu_{1}\rightarrow\infty$ by (181).
In this regime, we can approximate the water levels in (181) as follows
@@ H-B Low-Distortion Compression
$=$ $\frac{1}{2\nu_{1}}-\frac{1}{8\nu_{1}^{2}\lambda_{\ell}}+O\left(\frac{1}{\nu_{1}^{3}}\right).\;\;\;$
@@
$=$ $\frac{L}{2\nu_{1}}-\frac{1}{16\nu_{1}^{2}}\sum_{\ell=1}^{L}\frac{1}{\lambda_{\ell}}+O\left(\frac{1}{\nu_{1}^{3}}\right),$
@@
Substituting (191) into (189) shows that the water levels in the low-distortion regime are given by
$\gamma_{\ell}^{*}(\epsilon,0)$ $=$ $\frac{\epsilon}{L}-\frac{\epsilon^{2}}{2L^{2}\lambda_{\ell}}+\frac{\epsilon^{2}}{4L^{3}}\sum_{\ell=1}^{L}\frac{1}{\lambda_{\ell}}+O(\epsilon^{3}),$ (192)
$\hskip 85.35826pt\ell\in\{1,\ldots,L\}.$
@@
$R(\epsilon,0)$ (195)
$=$ $\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{\lambda_{\ell}}{\gamma^{*}_{\ell}(\epsilon,0)}$
$=$ $\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{L\lambda_{\ell}}{\epsilon}$
$-\frac{1}{2}\sum\limits_{\ell=1}^{L}\log\Bigg{(}1-\frac{\epsilon}{2L\lambda_{\ell}}+\frac{\epsilon}{4L^{2}}\sum_{\ell^{\prime}=1}^{L}\frac{1}{\lambda_{\ell^{\prime}}}+O(\epsilon^{2})\Bigg{)}\quad\;\;\;$
$=$ $\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{L\lambda_{\ell}}{\epsilon}-\frac{1}{2}\sum\limits_{\ell=1}^{L}\left(-\frac{\epsilon}{2L\lambda_{\ell}}+\frac{\epsilon}{4L^{2}}\sum_{\ell^{\prime}=1}^{L}\frac{1}{\lambda_{\ell^{\prime}}}\right)$
$\hskip 156.49014pt+O(\epsilon^{2})$